Given that A=3ax+5ay-7az and B=ax-2ay+az find
1
Expert's answer
2021-09-07T00:33:17-0400

"2B + 0.4A = <12*1+0.4*3,12*(-2)+0.4*5 , 12*1+0.4*(-7)>\\
 =<13.2, -22, 9.2> = 13.2 ax-22ay + 9.2az\\
A.B = 3*1 + 5*(-2) + (-7)*1 = 3-10-7 =-14\\
|B|^2 = 12 + (-2)^2 +12 = 1+4+1 = 6\\
A.B -|B|^2 = -14 - 6 =-20\\"